# Reconstructing MSM Sexual Networks to Guide PrEP Distribution Strategies for HIV Prevention

### v1

@@ Discussion
We developed a network reconstruction framework that generates realistic MSM sexual networks from individual-level behavioral and sociodemographic data, typically collected in HIV surveillance studies. While previous approaches have generated networks with enhanced clustering or degree correlations [2, 46, 7], ours is the first to systematically incorporate assortativity by age and race alongside these properties, capturing the demographic homophily that commonly shapes HIV transmission in MSM populations [1, 3]. This addresses a critical limitation: most HIV modeling studies either require complete contact data, extremely difficult to obtain for stigmatized populations, or use oversimplified network models that overlook key transmission dynamics.
Network topology significantly influenced both baseline HIV dynamics and intervention effectiveness. Assortative by degree networks showed the lowest equilibrium prevalence and fastest convergence, consistent with prior findings [14, 4]. Effectively, the increased fragmentation and longer path lengths in these networks create natural barriers to transmission. However, our study highlighted that community structure, not global metrics, emerged as the main determinant of intervention strategies efficiency. In clustered and assortative by degree networks, hubs concentrated within specific high-degree communities, creating transmission reservoirs where infection circulates rapidly among highly connected individuals [33, 52] but spreads slowly to the broader, lower-degree population. Consequently, for moderate adherence levels (80%) targeted PrEP strategies required higher coverage thresholds to outperform random allocation in these networks. Conversely, in assortative by age and race networks, hubs are distributed across communities and connect diverse demographic groups, positioning them as critical transmission bridges. Here, targeted strategies achieved maximal impact at lower coverage levels. For high adherence levels (95%) degree and position based allocation reduced prevalence substantially at only 40% coverage, while random distribution required double this coverage for equivalent effects. These findings are extremely valuable, validating theoretical predictions that identifying influential spreaders maximizes intervention efficiency in heterogeneous networks [38, 5], and enabling public health programs to achieve substantial epidemic control with fewer resources, lower costs, and reduced implementation complexity.
@@
While coverage and optimal PrEP allocation strategies are crucial, adherence proved equally critical to intervention success. At low adherence (23%), all strategies achieved minimal impact regardless of coverage or targeting approach, underscoring that efficient allocation alone cannot compensate for poor adherence. These findings stand out how crucial health education programs, PrEP counseling, and peer support are to reduce stigma, discrimination, and misconceptions about PrEP safety and efficacy [26], ensuring that network-optimized strategies achieve their full potential.
To assess real-world deployment, we compared empirical PrEP uptake patterns against simulated strategies, revealing substantial optimization opportunities. Current empirical distribution approximates optimal strategies in assortative by degree networks, where coverage exceeds 20% among high-degree individuals. However, in assortative by age and race networks, where hubs bridge demographic groups, both degree and $k$-shell targeting substantially outperformed empirical data, in which a lower PrEP coverage is present within demographic groups. This suggests that current eligibility criteria, focused primarily on individual risk behaviors, may fail to identify network-critical individuals whose PrEP initiation and uptake would maximally disrupt transmission chains. This strongly encourages PrEP allocation programs to incorporate network position, i.e., contact patterns, alongside behavioral risk factors, prioritizing individuals who bridge demographic groups or occupy central network positions.
@@
Our framework presents some limitations. Network generation without connectivity constraints resulted in the formation of a giant component and several disconnected components. This fragmentation was most pronounced in assortative by degree networks, where 16% of nodes were excluded from the largest component. Nevertheless, real-world sexual networks exhibit similar fragmentation patterns, supporting our modeling approach. Additionally, imposing connectivity would prevent achieving the high clustering and assortativity coefficients observed empirically [7], reflecting a fundamental trade-off between topological fidelity and completeness.
@@
Network data remain prohibitively difficult to collect in stigmatized populations. Without real-world contact patterns available, we could not validate our synthetic MSM sexual networks against empirical network structures, limiting assessment of our model’s generalizability. Additionally, most HIV surveillance studies present cross-sectional designs, which limits mathematical modelling of intervention strategies to accommodate dynamic features such as partnership formation and dissolution, shifting concurrency patterns, and potential risk compensation following PrEP adoption.
@@
Our work provides a practical methodology for optimizing PrEP distribution when complete network data are unavailable, highly common in HIV surveillance programs. Our findings demonstrate that network-informed allocation strategies could substantially improve HIV prevention efficiency in MSM populations, particularly in networks where hubs bridge demographic groups. These results establish benchmarks for future longitudinal network studies to validate synthetic network approaches and real-world effectiveness.
@@ B. Network Reconstruction of Null Models to represent Empirical Networks
We consider a graph $G=(V,E)$, as an ordered tuple pair of finite sets, through which we aimed to represent our dataset. The elements of $V$ represent the nodes of the network, which correspond to the study participants, connected by edges $e\in E\subseteq V\times V$, the sexual interactions between them. The resulting network obtained from the data preprocessing is composed of $V=4667$ nodes and $E=25107$ edges, such that each node is characterized by the following set of attributes: age, race, number of partnerships, testing frequency and PrEP status.
Since the network structure was absent, to generate the networks, we used the number of partnerships reported by each participant and assume they are connected in the real world. In the context of generating random graphs with a prescribed degree sequence, let $d_{i}$ represent the degree sequence of graph $G$. Since the sum of all degrees in $G$ denoted by $\sum_{i=0}^{4667}d_{i}=50214$ is even, random graphs can be constructed pairing up all nodes. We developed an algorithm based on the Havel-Hakimi [24, 23] recursive test that provides a framework to build a simple, undirected, connected graph meeting the above degree sequence, when the conditions postulated by Erdős-Gallai are obeyed [18, 10]. This algorithm was preferred over the classical configuration model since it does not generate self-loops and duplicate edges [8], which are not desirable to represent sexual networks. In our approach, nodes are ordered in descending order according to their degree, such that higher-degree nodes are the first ones to get paired up. The algorithm stops when the stubs of each node are connected (see Supplementary Information Algorithm 1 and Section 2).
Despite its advantages in the generation of simple graphs, the Havel-Hakimi algorithm is greedy and generates a highly correlated graph configuration space. Therefore, the capability of generalizing the macroscopic structure of our dataset can be compromised. To address this limitation, a MCMC approach was implemented using the method double edge swap from NetworkX. Let $G_{t}$ be the graph configuration at time $t$. With probability 0.5, $G_{t+1}$ is determined through an edge swapping process. We pick two edges at random $(u,v)$ and $(x,y)$, with distinct endpoints. If $(u,x)$ and $(v,y)$, are not connected, we remove the old edges and replace them by the proposed ones. To maximize exploration of the configuration space, no connectivity constraint was imposed during edge swapping. Since no prior estimates of mixing time exist for such networks, we ran chains until convergence, assessed by evaluating the convergence of the clustering coefficient, degree assortativity, and diameter. The pseudo-code of the algorithm and detailed convergence analysis can be found in Supplementary Information, Fig. S4, Algorithm 2 and Section 3.
@@ C. Epidemiological and Agent Based Model
An agent-based model was developed to model HIV disease spreading in the networks with the highest enhanced clustered and assortative networks by degree, age and race to mimic real-world systems. The model is assumed to be static, which implies that the degree and the nodes each agent interacts with remain constant throughout the entire simulation. The time evolution of the nodes status is modeled by a Markov process. A rejection sampling algorithm was developed to simulate the stochastic process of the disease spreading. Therefore, time was assumed to be discrete and fixed for a $\Delta t=1$ year. Since sexual networks are formed of multiple disconnected components, we did not impose constraints on the network’s connectivity in the network generation process. Therefore, the simulation model was applied in the largest component only.
Considering a population of size $N$, each individual is in one of three states: Susceptible $S$, Infected $I$, and Treatment $T$, which corresponds to individuals who start to take Antiretroviral Therapy (ART) after testing positive for HIV. Individuals who are on ART are assumed to reach a zero viral load, in which individuals are not able to transmit the disease anymore [13]. To simulate the disease transmission we considered $t_{\text{end}}$ = 50 years and $N_{\text{trials}}$ = 50, to reduce bias and assess the robustness of the model to randomness. We chose to implement the stochastic epidemic process in a synchronous way, such that the state of each node is only updated after all nodes have been traversed. Through this approach, each node is given the possibility to transition to another infectious stage. A Bernoulli trial was performed per node to determine if the node will change state. The transition between states will be determined by two main parameters, $\beta$, the probability of a susceptible node to become infected and $\gamma$ the probability of an infected node getting tested and initiating treatment. If the current node is susceptible and presents infected neighbors, such that $u<\beta$, the node transitions from $S\rightarrow I$. There is, however, two types of susceptible individuals in our model, on PrEP and not on PrEP, such that nodes on PrEP, have a lower probability of contracting HIV. Additionally, the infection probability in general also depends on the number of neighbor and the type of interactions, which includes the usage of other prevention methods, such as condoms, and the type of sexual role. The impact of these factors on transition probabilities are discussed in detail below. Finally, if the current node is infected and $u<\gamma$, the node transitions from $I\rightarrow T$. The pseudo-code of this algorithm is provided in the Supplementary Information Algorithm 3.
@@ D. Parameter Estimation
The testing frequency of each agent retrieved from the dataset will determine the probability of getting tested, expressed as
@@
The higher the number of times a person gets tested per year, the higher the likelihood of getting tested. Additionally, we multiply the probability by a 1/2 factor, since each agent reported cumulatively its testing frequency for 24 months, when we conduct simulations with a time increment of $\Delta t=1$ year.
@@
In order to take into account the heterogeneity of each node’s neighbourhood in HIV transmission, a binomial model was defined according to Newman et. al [34]. The overall probability of infection $\beta$ will depend on the probability a susceptible node escapes from getting infected, which is constant, and on the number of their infected neighbours $n$. Assuming that time $t$ is discrete we have that:
@@ E. Targeted PrEP allocation Strategies
The efficiency of an immunization strategy also depends on the network properties [38]. Therefore, we compared the effect of choosing nodes uniformly at random with two targeting strategies, based on the node’s degree centrality and the k-shell decomposition method. In the degree centrality approach, nodes are prioritized based on their degree, while in the k-shell method, prioritization depends on how nodes are connected and their location in the network. We start by decomposing the network into $k$-shells. First, we remove all the nodes that present only one link, until no more nodes with a remaining degree equal to 1 are present in the network, and we add them to a sub-list which we will denote 11-shell. Next, we remove all the nodes that have degree 2 and add them to a sub-list called 22-shell. We continue increasing $k$ and to apply this procedure until all nodes were attributed to a $k$- shell. The $k$-core corresponds to all $k$- shells with indices higher or equal to $k$. Nodes that belong to the shell with the highest $k$ form the nucleus [11]. These approaches were chosen since, when compared with other centrality measures, they are the most efficient to identify influential nodes in the context of epidemics [17], especially when the epidemic starts in multiple origins simultaneously [28].
@@ Appendix 1 Figures
Figure S4: Markov Chain convergence obtained for three different networks generated through the Havel-Hakimi algorithm. The MCMC approach was applied eight times for each starting network during 200,000 steps. The median and respective interquartile range of successful edge swappings were 91030.0 (90883.25-91181.5). The figure presents the median and interquartile variation of the transitivity (a), assortativity by degree coefficient (b) and the sampled diameter (c) throughout the simulation.
Figure S5: Network evolution as we increase clustering, assortativity by degree, age and race. The network generation algorithm was simulated for 500,000 steps such that, most networks converge after, approximately 200,000 steps.
Figure S6: Similarity metrics of clustering methods. Normalized Mutual Information (NMI) (a) and Variation of Information (VI) (b) considering N = 300 stochastic trials, obtained for clustered, assortative by degree, race, and age, and considering different partition methods, the Louvain algorithm and the NSBM.
(a) Evolution of structural metrics (diameter, average shortest path length, largest component size, number of disconnected components, and modularity) as clustering and assortativity coefficients increase over 500,000 steps.
(b) HIV prevalence over 50 years across networks with different topologies. Each line shows the median and interquartile range from $N=30$ stochastic simulations.

### v2

@@ Discussion
We developed a network reconstruction framework that generates realistic MSM sexual networks from individual-level behavioral and sociodemographic data, typically collected in HIV surveillance studies. While previous approaches have generated networks with enhanced clustering or degree correlations [2, 50, 8], our study is the first to systematically incorporate and evaluate assortativity by age and race alongside these properties, capturing the demographic homophily that commonly shapes HIV transmission in MSM populations [1, 3]. This addresses a critical limitation: most HIV modeling studies either require complete contact data, extremely difficult to obtain for stigmatized populations, or use oversimplified network models that overlook key transmission dynamics.
We show that network topology significantly influences both baseline HIV dynamics and intervention effectiveness. Assortative by degree networks showed the lowest equilibrium prevalence and fastest convergence, consistent with prior findings [16, 4]. Indeed, the increased fragmentation and longer path lengths in these networks create natural barriers to transmission. However, our study highlights that community structure, not global metrics, emerge as the main determinant of intervention strategies efficiency. In clustered and assortative by degree networks, hubs are concentrated within specific high-degree communities, creating transmission reservoirs where infection circulates rapidly among highly connected individuals [36, 56] but spreads slowly to the broader, lower-degree population. Consequently, for moderate adherence levels (80%), targeted PrEP strategies require higher coverage thresholds to outperform random allocation in these networks. Conversely, in assortative by age and race networks, hubs are distributed across communities and connect diverse demographic groups, positioning them as critical transmission bridges. Here, targeted strategies achieve maximal impact at lower coverage levels. For high adherence levels (95%) degree and position based allocation reduced prevalence substantially at only 40% coverage, while random distribution required double this coverage for equivalent effects. These findings are extremely valuable, validating theoretical predictions that identifying influential spreaders maximizes intervention efficiency in heterogeneous networks [41, 5]. This can inform public health programs to achieve substantial epidemic control with fewer resources, lower costs, and reduced implementation complexity.
@@
While coverage and optimal PrEP allocation strategies are crucial, adherence is equally critical to intervention success. At low adherence (23%), all strategies achieved minimal impact regardless of coverage or targeting approach, underscoring that efficient allocation alone cannot compensate for poor adherence. These findings show how crucial health education programs, PrEP counseling, and peer support can be to reduce stigma, discrimination, and misconceptions about PrEP safety and efficacy [29]. This could motivate a higher adherence and facilitate network-optimized strategies to achieve their full potential.
To assess real-world implementation, we compared PrEP uptake retrieved from the data with simulated strategies, revealing substantial optimization opportunities. When using information about who is on PrEP for network reconstruction, its uptake in the largest component of each network type approximates optimal strategies in assortative-by-degree networks, where coverage exceeds 20% among high-degree individuals. However, in assortative-by-age and assortative-by-race networks, where hubs bridge demographic groups, both degree and $k$-shell targeting substantially outperformed the empirical distribution, in which lower PrEP coverage is observed within these demographic hubs. Although eligibility criteria typically only consider individual risk behaviors, our results suggest that incorporating network-critical information could significantly improve the identification of individuals whose PrEP initiation and uptake would maximally disrupt transmission chains.To contribute to this, future surveys could go beyond asking about the number of contacts by also collecting data on mixing preferences, which may translate into more effective PrEP allocation strategies. Crucially, because we cannot determine from egocentric data alone which assortativity dimension best reflects the ground-truth network, our findings underscore structural uncertainty as a primary challenge for PrEP efficacy. By clarifying the interplay between network structure and PrEP allocation strategies, this work highlights the importance of investing in richer network data collection to support allocation strategies that remain robust under structural uncertainty.
@@
Our framework presents, however, some limitations. Network generation without connectivity constraints resulted in the formation of a giant component and several disconnected components. This fragmentation was most pronounced in assortative by degree networks, where 16% of nodes were excluded from the largest component. Nevertheless, since real-world sexual networks exhibit similar fragmentation patterns [27], we believe that our modeling approach is still valid. Additionally, imposing connectivity would prevent achieving the high clustering and assortativity coefficients observed empirically [8], reflecting a fundamental trade-off between topological fidelity and completeness. A key direction for future work is to jointly analyze multiple network properties—such as degree assortativity, clustering, and demographic assortativity by age and race—since in real sexual networks these features are often correlated rather than independent.
@@
Network data remain prohibitively difficult to collect in stigmatized populations. Without real-world contact patterns available, we could not validate our synthetic MSM sexual networks against empirical network structures, limiting assessment of our model’s generalizability. Additionally, several HIV surveillance studies present cross-sectional designs [12], which limits mathematical modelling of intervention strategies to accommodate dynamic features such as partnership formation and dissolution, shifting concurrency patterns, and potential risk compensation following PrEP adoption.
@@
Overall, our work provides a practical methodology for the study of PrEP distribution optimization when complete network data are unavailable, highly common in HIV surveillance programs. Our findings demonstrate that network-informed allocation strategies could substantially improve HIV prevention efficiency in MSM populations, particularly in networks where hubs bridge demographic groups. These results establish benchmarks for future longitudinal network studies to validate synthetic network approaches and real-world effectiveness.
@@ B. Network Reconstruction of Null Models to represent Empirical Networks
We consider a graph $G=(V,E)$, in which the elements of $V$ represent the nodes of the network, representing the study participants, connected by edges $e\in E\subseteq V\times V$, representing the sexual interactions between them. The resulting graph (network) obtained from the data preprocessing is composed of $V=4667$ nodes and $E=25107$ edges, such that each node is characterized by the following set of attributes: age, race, number of partnerships, testing frequency and PrEP status.
Since the network structure is unknown, to generate the networks we used the number of partnerships reported by each participant and assume they are connected in the real world. In the context of generating random graphs with a prescribed degree sequence, let $d_{i}$ represent the degree sequence of graph $G$. Since the sum of all degrees in $G$ denoted by $\sum_{i=0}^{4667}d_{i}=50214$ is even, random graphs can be constructed pairing up all nodes. We developed an algorithm based on the Havel-Hakimi [26, 25] recursive test that provides a framework to build a simple, undirected, connected graph meeting the above degree sequence, when the conditions postulated by Erdős-Gallai are obeyed [20, 11]. This algorithm was preferred over the classical configuration model since it does not generate self-loops and duplicate edges [9], which are not desirable to represent sexual networks. In our approach, nodes are ordered in descending order according to their degree, such that higher-degree nodes are the first ones to get paired up. The algorithm stops when the stubs of each node are connected (see Supplementary Information Algorithm 1 and Section 2).
Despite its advantages in the generation of simple graphs, the Havel-Hakimi algorithm is greedy and generates a highly correlated graph configuration space. Therefore, the capability of generalizing the macroscopic structure of our dataset could be compromised. To address this limitation, a MCMC approach was implemented using the method double edge swap from NetworkX. Let $G_{t}$ be the graph configuration at time $t$. With probability 0.5, $G_{t+1}$ is determined through an edge swapping process. We select two edges at random $(u,v)$ and $(x,y)$, with distinct endpoints. If $(u,x)$ and $(v,y)$, are not connected, we remove the old edges and replace them by the proposed ones. To maximize exploration of the configuration space, no connectivity constraint was imposed during edge swapping. Since no prior estimates of mixing time exist for such networks, we ran chains until convergence, assessed by evaluating the convergence of the clustering coefficient, degree assortativity, and diameter. The pseudo-code of the algorithm and detailed convergence analysis can be found in Supplementary Information, Fig. S4, Algorithm 2 and Section 3.
@@ C. Epidemiological and Agent Based Model
We developed an agent-based model to model HIV disease spreading in the networks with the enhanced clustered and assortative networks by degree, age and race to mimic real-world systems. The model is assumed to be static, which implies that the degree and the nodes each agent interacts with remain constant throughout the entire simulation. The time evolution of the nodes status is modeled by a Markov process. A rejection sampling algorithm was developed to simulate the stochastic process of the disease spreading. Therefore, time was assumed to be discrete and fixed for a $\Delta t=1$ year. Since sexual networks are formed of multiple disconnected components, we did not impose constraints on the network’s connectivity in the network generation process. Therefore, the simulation model was applied on the largest component only.
Considering a population of size $N$, each individual is in one of three states: Susceptible $S$, Infected $I$, and Treatment $T$, which corresponds to individuals who start to take Antiretroviral Therapy (ART) after testing positive for HIV. Individuals who are on ART are assumed to reach a zero viral load, in which individuals are not able to transmit the HIV anymore [15]. To simulate the HIV transmission we considered $t_{\text{end}}$ = 50 years and $N_{\text{trials}}$ = 50, to reduce bias and assess the robustness of the model to randomness. We chose to implement the stochastic epidemic process in a synchronous way, such that the state of each node is only updated after all nodes have been traversed. Through this approach, each node is given the possibility to transition to another infectious stage. A Bernoulli trial was performed per node to determine if the node will change state. The transition between states will be determined by two main parameters, $\beta$, the probability of a susceptible node to become infected and $\gamma$ the probability of an infected node getting tested and initiating treatment. If the current node is susceptible and presents infected neighbors, such that $u<\beta$, the node transitions from $S\rightarrow I$. There is, however, two types of susceptible individuals in our model, on PrEP and not on PrEP, such that nodes on PrEP, have a lower probability of contracting HIV. Additionally, the infection probability in general also depends on the number of neighbor and the type of interactions, which includes the usage of other prevention methods, such as condoms, and the type of sexual role. The impact of these factors on transition probabilities are discussed in detail below. Finally, if the current node is infected and $u<\gamma$, the node transitions from $I\rightarrow T$. The pseudo-code of this algorithm is provided in the Supplementary Information Algorithm 3.
@@ D. Parameter Estimation
The testing frequency of each agent retrieved from the data determines the probability of getting tested, expressed as
@@
The higher the number of times a person gets tested per year, the higher the likelihood of getting tested. Additionally, since we run simulations with a time increment of $\Delta t=1$ year, we multiply this probability by a 1/2 factor, since each agent reported cumulatively its testing frequency for 24 months.
@@
In order to take into account the heterogeneity of each node’s neighbourhood in HIV transmission, a binomial model was defined according to Newman et. al [37]. The overall probability of infection $\beta$ depends on the probability that a susceptible node escapes from getting infected, which is constant, and on the number of their infected neighbours $n$. Assuming that time $t$ is discrete we have that:
@@ E. Targeted PrEP allocation Strategies
The success of an immunization strategy also depends on the network properties [41]. Therefore, we compared the effect of choosing nodes uniformly at random with two targeting strategies, based on the node’s degree centrality and the k-shell decomposition method. In the degree centrality approach, nodes are prioritized based on their degree, while in the k-shell method, prioritization depends on how nodes are connected and their location in the network. We start by decomposing the network into $k$-shells. First, we remove all the nodes that present only one link, until no more nodes with a remaining degree equal to 1 are present in the network, and we add them to a sub-list which we will denote 11-shell. Next, we remove all the nodes that have degree 2 and add them to a sub-list called 22-shell. We continue increasing $k$ and to apply this procedure until all nodes were attributed to a $k$- shell. The $k$-core corresponds to all $k$- shells with indices higher or equal to $k$. Nodes that belong to the shell with the highest $k$ form the nucleus [13]. These approaches were chosen since, when compared with other centrality measures, they are the most efficient to identify influential nodes in the context of epidemics [19], especially when the epidemic starts in multiple origins simultaneously [31].
@@ Acknowledgments
Appendix 1 Section 1: KNN Imputation
Appendix 2 Section 2: Havel-Hakimi
The proposed Havel-Hakimi algorithm receives as an input a list $D$ of tuples, such that each tuple corresponds to the ID of the current node and its respective degree. Per iteration, $D$ is ordered in descending order such that the tuple that presents the highest degree is removed from $D$. This node will be denoted by ego. When the ego is removed, the resultant list $D^{\prime}$ will be composed by the remaining nodes and respective degrees. After, the selected node will be connected sequentially to the other nodes, such that the nodes that present a higher degree have priority. This explains why the Havel-Hakimi algorithm presents itself as a greedy algorithm. As the current ego is connected to each node from $D^{\prime}$, the number of available stubs is decreased such that their degrees are updated, respectively, to D[0][1]1D[0][1]-1 and $D^{\prime}[i][1]-1$, where $i$ corresponds to the current node from $D^{\prime}$ selected to be connected with the ego. When all stubs from the current ego are connected ($D[0][1]=0$), the current ego is updated and becomes the first node from $D^{\prime}$. For the degree sequence to be graphical, the degrees of successive $D^{\prime}$ sequences obtained by removing the ego from $D$ should always be superior or equal to zero. Otherwise, the number of edges of the degree sequences changes for the algorithm to produce a graphical sequence, which explains why, in the final step, we verify if the number of edges of the generated graph with the prescribed degree sequence $L$ remains equal to the number of initial edges. We generated 10 networks through the Havel-Hakimi algorithm.
Appendix 3 Section 3: MCMC Convergence Analysis
To assess the convergence of the uniform sampling following a MCMC approach, we measured the transitivity, assortativity by degree, and the sampled diameter [23] (see Figure S4) of 100 random networks, sampled from 10 initial networks obtained from the Havel-hakimi algorithm. Initially, due to the greedy nature of the Havel-Hakimi, networks present high transitivity, assortativity by degree, and diameters. However, as the number of successful edge swaps increased, the networks reached a stationary state characterized by vanishing clustering and assortativity coefficients, as well as reduced diameters, indicating a transition towards random mixing.
Appendix 5 Section 5: Community Detection
We applied the Nested Stochastic Block Model (NSBM), a probabilistic approach to address the reported limitations of modularity maximization algorithms. For each stochastic realization of the 100 clustered, assortative by degree, age and race networks, we obtained $N=50$ partitions using the NSBM. For each stochastic trial and network type, we chose the partition that registered the lowest descriptive length [44]. We evaluated the robustness of the partitions obtained for each network stochastic trial by obtaining the distribution of the number of communities detected, size of the largest community and average degree (see Figure S7), which was very similar across network types for different partitions obtained for each stochastic trial. To quantify this consistency, we applied two-sample Kolmogorov–Smirnov tests to community size distributions from 100 realizations per network type (Figure S8). Under the null hypothesis that pairwise community size distributions come from the same distribution, most comparisons within each network type failed to reject $H_{0}$ ($p>0.05$): 83.9% for clustered networks, 99.4% for degree-assortative, 99.88% for age-assortative, and 100% for race-assortative networks.
@@ Appendix 6 Figures
Figure S4: Markov Chain convergence obtained for ten different networks generated through the Havel-Hakimi algorithm. We generated 10 $MC$ networks for each Havel-Hakimi network during 200,000 steps. The median and respective interquartile range of successful edge swappings were 91030.0 (90883.25-91181.5). The figure presents the median and interquartile variation of the transitivity (a), assortativity by degree coefficient (b) and the sampled diameter (c) throughout the simulation.
Figure S5: Network properties coefficient evolution during network reconstruction based on clustered and assortativity by degree and sociodemographic attributes. The figure presents the median and interquartile variation of 100 clustered, assortativity by degree, age and race networks, simulated for 500,000 steps.
(a) Evolution of structural metrics (diameter, average shortest path length, largest component size, number of disconnected components, and modularity) as clustering and assortativity coefficients increase over 500,000 steps. Each global property was represented through their respective mean and standard deviation over 100 MCMC networks obtained for each network type.
(b) HIV prevalence over 50 years across networks with different topologies. Each line shows the median and interquartile range from $N=50$ stochastic simulations.
@@
Figure S7: Structural characterization of community partition obtained by the Nested Stochastic Block Model over 100 networks realizations for each network type. We present, for each network type, the median and interquartile range of the number of communities (a), largest community size (b) and average degree (c).
Figure S8: Two-Sample Kolmogorov–Smirnov Test on the community size distribution for the partitions obtained for the 100 realizations of each network type. We present the p-values obtained for each pairwise combination of each distribution of number of nodes across partitions for clustered (a), assortativity by degree (b), age (c) and race (d) networks. The dashline represents the threshold below which the null hypothesis $H_{0}$ that any two distributions come from the same distribution should be rejected.